OpenStudy (anonymous):

solve for y x=1/8(y-4)

OpenStudy (anonymous):

\[y-4 = \frac{1}{8x}\] \[y = \frac{1}{8x}+4\]

OpenStudy (anonymous):

1/8x+4?

OpenStudy (lalaly):

\[8x=y-4\]\[y=8x+4\]

OpenStudy (lalaly):

its \[x=\frac{1}{8}(y-4)\]right?

OpenStudy (anonymous):

yeah

OpenStudy (lalaly):

then its \[y=8x+4\]

OpenStudy (anonymous):